# Reducing base drag on road vehicles using pulsed jets optimized by hybrid genetic algorithms

### v1

@@ Abstract
Aerodynamic drag on flat-backed vehicles like vans and trucks is dominated by a low-pressure wake, whose control is critical for reducing fuel consumption. This paper presents an experimental study at $Re_{W}\approx 78,300$ on active flow control using four pulsed jets at the rear edges of a bluff body model. A hybrid genetic algorithm, combining a global search with a local gradient-based optimizer, was used to determine the optimal jet actuation parameters in an experiment-in-the-loop setup. The cost function was designed to achieve a net energy saving by simultaneously minimizing aerodynamic drag and penalizing the actuation’s energy consumption. The optimization campaign successfully identified a control strategy that yields a drag reduction of approximately 10%. The optimal control law features a strong, low-frequency actuation from the bottom jet, which targets the main vortex shedding, while the top and lateral jets address higher-frequency, less energetic phenomena. Particle Image Velocimetry analysis reveals a significant upward shift and stabilization of the wake, leading to substantial pressure recovery on the model’s lower base. Ultimately, this work demonstrates that a model-free optimization approach can successfully identify non-intuitive, multi-faceted actuation strategies that yield significant and energetically efficient drag reduction.
@@
Nomenclature
  • $\alpha$
    Non-dimensional reduced coordinates
  • $U_{\infty}$
    Free-stream velocity
  • $V_{jet}$
    Mini-jet exit velocity
  • $V_{suction}$
    Boundary layer suction velocity
  • $W$
    Width of the experimental model
  • $H$
    Height of the experimental model
  • $L$
    Length of the experimental model
  • $h$
    Ground clearance of the experimental model
  • $L_{B}$
    Length of the primary moment arm of the cantilever
  • $L_{b}$
    Length of the secondary moment arm of the cantilever
  • $L_{j}$
    Distance between the Ahmed body and the boundary layer removal jet centerline
  • $P_{j}$
    Actuator upstream pressure
  • $\Delta P$
    Base pressure differential
  • $f$
    Frequency
  • DCDC
    Duty Cycle
  • $\phi$
    Phase angle
  • $\tilde{\bullet}$
    Non-dimensional operator
  • $\overline{\bullet}$
    Time-average operator
  • $\bullet^{\prime}$
    Fluctuating operator
  • $Re_{W}$
    Width-Reynolds Number
  • $K$
    Control Law
  • $J$
    Cost function value
  • $J_{a}$
    Drag-cost value
  • $J_{b}$
    Penalization value
  • $D$
    Aerodynamic Drag
  • $\gamma$
    Penalization constant
  • x,y,zx,\;y,\;z
    Stream-wise, span-wise, and wall-normal coordinates
  • U,VU,\;V
    Stream-wise and wall-normal velocity components
  • $p$
    Wall pressure
  • $\omega$
    Vorticity
  • $T_{w,D_{0}}$
    Waiting time for the non-actuated section
  • $T_{m,D_{0}}$
    Measurement time for the non-actuated section
  • $T_{w,D}$
    Waiting time for the actuated section
  • $T_{m,D}$
    Measurement time for the actuated section
@@ 1 Introduction
Road freight is responsible for approximately 5.2% of global greenhouse-gas emissions [1]. At highway speeds, roughly 65% of a vehicle’s tractive energy is expended to overcome aerodynamic drag [2]. For flat-backed vehicles such as vans and lorries, this drag is dominated by a low-pressure region on the base, which results from flow separation at the sharp trailing edges. Consequently, mitigating this base-pressure deficit through wake control provides a direct and effective means to reduce fuel consumption and the associated emissions.
@@
The wakes of simplified ground vehicles like the Ahmed body have been the subject of extensive research, revealing a highly complex, three-dimensional flow field that is acutely sensitive to geometric variations [3]. The near-wake is typically characterized by two large recirculation regions originating from the shear layers that separate from the upper and lower base edges [4]. In the lateral direction, the wake often exhibits a global symmetry breaking, with modal switching between quasi-symmetric and anti-symmetric states. This low-frequency dynamic, linked to the shedding of large-scale vortex structures, produces broadband pressure fluctuations and a depressed time-averaged base pressure. Foundational experimental and numerical studies have thoroughly mapped the bi-modal character of these wakes and their sensitivity to geometry and operating conditions [5, 6, 7], providing a baseline understanding that informs all modern control strategies.
@@
Modifying these wake dynamics offers a promising route for drag reduction. Passive control devices have been widely explored, including vertical flaps to attenuate lateral vortex structures [8], base cavities to improve performance in crosswinds [9], splitter plates [10], boat tails [11, 12], and fences [13]. In contrast, active flow control provides greater authority due to its adaptive capabilities to changing conditions [14] and has been investigated extensively through momentum injection at the trailing edges [15]. In particular, strategies involving steady momentum injection have proven effective, with Rouméas et al. [16] achieving a 17% drag reduction on an Ahmed body via suction prior to separation, and Aubrun et al. [17] obtaining up to 14% drag reduction using an array of steady blowing microjets. A widespread and often more efficient alternative is unsteady forcing, through pulsed or synthetic jets. These actuators can promote flow reattachment, as explored by Glezer et al. [18], or directly target wake instabilities. Examples of the latter include parametric studies to optimize actuation on the slanted surfaces of an Ahmed body [19, 20], and targeted base pressure recuperation using high-frequency forcing [21]. Some approaches have even combined passive devices with active suction and blowing to maximize performance [22].
@@
More recently, learning-based, model-free approaches have gained traction, offering powerful tools to navigate the vast and complex parameter space of active flow control laws directly on experimental setups. This paradigm bypasses the need for explicit, often intractable, low-dimensional models of turbulent wakes, allowing algorithms to discover effective, sometimes non-intuitive, control strategies through interaction with the physical system [23, 24, 25]. Several successful applications have demonstrated the potential of these methods on benchmark automotive geometries, primarily the Ahmed body. For instance, Li et al. [23] employed linear genetic programming (LGP) to optimize pulsed jets combined with Coanda surfaces on a square-back Ahmed body, achieving a remarkable 22% drag reduction. Their model-free approach identified optimal multi-frequency forcing laws. Zhang et al. [24] utilized an Ant Colony Optimization (ACO) algorithm to control distributed steady microjet arrays on a low-drag Ahmed body (slant angle $\varphi=35^{\circ}$), reducing drag by 18% while explicitly incorporating control power input into the cost function to seek efficient solutions. Deng et al. [26] applied an Explorative Downhill Simplex Method (EDSM) to optimize independently operated pulsed microjets on a square-back Ahmed body, achieving 13% drag reduction. Their work particularly focused on sensitivity analysis for a large number of control parameters (up to 12) and demonstrated the potential for significantly enhanced control efficiency (up to 78%) with only a minor sacrifice in drag reduction. Demonstrating a fully autonomous approach, Zhang et al. [25] developed REACT, a Reinforcement Learning framework deployed experimentally on a square-back Ahmed body using servo-actuated flaps. Learning directly in the wind tunnel from onboard pressure sensors, REACT achieved a 3.64% drag reduction with net energy savings, discovering that dynamically suppressing wake instabilities was the optimal strategy. Notably, their physics-informed training allowed the agent to generalize across different flow speeds without retraining. Collectively, these successes underscore the competitive performance and adaptability of machine learning control for manipulating complex wake flows on standard automotive shapes
@@
Moving beyond the standard Ahmed body benchmark, active flow control strategies have also been applied to more realistic vehicle shapes, notably the van model geometry utilized in [27, 28, 29]. In a first study, a Genetic Algorithm (GA) was used to optimize multi-frequency pulsed jets in a square-back van (represented as a sum of two sine waves), achieving an 11.2% drag reduction. Notably, their initial cost function aimed purely at maximizing drag reduction without explicitly penalizing actuation energy. The follow-up work [29] employed Deep Reinforcement Learning (DRL), exploring different state representations (using base pressure) and reward definitions, including cases explicitly designed to consider the energy budget alongside drag reduction. Their subsequent flow topology analysis revealed distinct wake modifications depending on whether the DRL agent prioritized maximal drag reduction or energy efficiency.
@@
Despite the demonstrated promise, the efficacy and practical applicability of these learning-based strategies can be constrained by several factors. Firstly, the reliance on a single optimization algorithm (be it GA, LGP, ACO, EDSM, or DRL) within many individual studies risks incomplete exploration of the potentially vast and multimodal control parameter space, possibly leading to convergence towards locally optimal, rather than globally optimal, solutions. Secondly, a critical consideration, often omitted in earlier works but highlighted in more recent ones, is the actuation cost. Neglecting the energy expenditure of the control system within the optimization loop can yield strategies that, while achieving substantial drag reduction, are energetically inefficient or entirely impractical for real-world applications. While several studies have begun incorporating energy considerations [e.g. 24, 26, 29], it remains a crucial factor for practical viability. Finally, many experimental implementations, particularly those involving hardware-in-the-loop optimization, do not formally account for the propagation of measurement uncertainty. The inherent noise in experimental data, if not properly managed, can be amplified, introducing significant bias into the learning process and potentially guiding the optimization towards spurious or non-robust outcomes.
@@
To address these limitations, the present study introduces a novel hybrid genetic algorithm that synergistically combines the global exploratory power of a genetic algorithm with the local exploitation capabilities of the Downhill Simplex Method, a local search algorithm. We apply this algorithm experimentally to a simplified van model in a wind tunnel as in Amico et al. [27, 28, 29], optimizing the control parameters of a pulsed-jet system while simultaneously characterizing the resulting wake modifications. Crucially, the objective function incorporates a penalty term for the momentum injected by the actuators, ensuring that the optimization converges towards energetically feasible solutions. Furthermore, to guarantee the robustness of the experimental evaluation, each candidate solution is subjected to repeated measurements, and a statistical uncertainty threshold is employed to discard unreliable data, thereby providing a high degree of confidence in the final performance metrics.
@@
The remainder of this paper is organized as follows. Section 2 details the experimental setup, including the wind tunnel facility, the bluff body model, and the actuation and measurement systems. Section 3 presents the hybrid genetic algorithm, defining the control parameterization, the cost function, and the methodology for handling experimental uncertainty. The results of the optimization are presented and discussed in Section 4, where the performance of the optimal control strategies is analyzed in detail, and the corresponding flow physics are examined in Section 4.3. Finally, Section 5 summarizes the key findings of the study and offers concluding remarks.
@@ 2.1 Experimental Facility and Model Setup
The experiments were conducted in an Eiffel-type wind tunnel at the Spanish National Institute for Aerospace Technology (INTA), Spain. The facility features a test section with a square cross-section of $0.4\times 0.4$ m2 and a length of $2$ m. All measurements were performed at a free-stream velocity of $U_{\infty}=12$ m/s, with a corresponding turbulence intensity below 0.3%.
The bluff body is a scaled version of the van model used in Cerutti et al. [4] and Amico et al. [27, 28, 29], with a height of $H=0.12$ m, a width of $W=0.102$ m, and a length of $L=0.247$ m. The clearance between the model and the wind tunnel floor is equal to $h=0.012$ m, leaving a normalized value $h/H=0.1$. This configuration results in a wind tunnel blockage ratio of 7.65% and a Reynolds number based on width of $Re_{W}\approx 78,300$. It is noted that, even though the external geometry of the model is identical to the referenced study, the interface with the load transducer and the actuation systems is not the same (see §2.2). As depicted in Figure 1, the model is positioned on a free-floating platform, maintaining a lateral clearance of approximately $1$ mm with respect to the wind tunnel floor. It is rigidly mounted to the platform through rectangular extensions from the wheels, designed so that the wheel base would be flush with the floor. This mounting strategy provides a non-intrusive passage for the tubing of the actuation and instrumentation systems, thereby precluding the need for internal support structures and the associated flow corrections.
@@ 2.2 Actuation System
The actuation system comprises four pairs of parallel slots positioned on the four edges of the model’s base, as indicated by the red lines in Figure 1, (b). This arrangement is analogous to actuator configurations used in recent bluff-body flow-control studies, such as the single-slot jets employed by Amico et al. [27]. Each rectangular slot features a cross-section of $39.4\times 1$ mm2. The use of slot pairs, separated by $5$ mm on each side of the model, is employed to improve the homogeneity of the jet outflow.
@@ 2.3 Force and base Pressure Measurements
Aerodynamic drag force is measured via a custom lever system acting on a Fibos FA702 three-axis load cell. The body-platform assembly, described in Section 2.1, is mounted on a primary cantilever beam of length $L_{B}=0.82$ m. This beam pivots on a high-precision bearing, a mechanism that translates the streamwise aerodynamic force exerted on the model into a vertical load on a secondary lever arm of length $L_{b}=85$  mm. The resulting force is transmitted via a vertical threaded rod to the load cell. In the primary measurement axis, the load cell has a maximum capacity of $5$ N and sensitivity 1mV/V1\;mV/V with an excitation of 12V12\;V. Data from the load cell is acquired using a Viking VK702NH data acquisition system, which provides a $10$ V excitation voltage and utilizes a 24-bit ADC with an input range of $100$ mV, sampling at $800$ Hz.

### v2

@@ Abstract
Aerodynamic drag on flat-backed vehicles like vans and trucks is dominated by a low-pressure wake, whose control is critical for reducing fuel consumption. This paper presents an experimental study at $Re_{W}\approx 78,300$ on active flow control using four pulsed jets at the rear edges of a bluff body model. A hybrid genetic algorithm, combining a global search with a local gradient-based optimizer, was used to determine the best-performing jet actuation parameters in an experiment-in-the-loop setup. The cost function was designed to achieve a net energy saving by simultaneously minimizing aerodynamic drag and penalizing the actuation’s energy consumption. The optimization campaign successfully identified a control strategy that yields a drag reduction of approximately 8.8%. The best-performing control law features a strong, low-frequency actuation from the bottom jet, which targets the main vortex shedding, while the top and lateral jets address higher-frequency, less energetic phenomena. Particle Image Velocimetry analysis reveals a significant upward shift and stabilization of the wake, leading to substantial pressure recovery on the model’s lower base. Ultimately, this work demonstrates that a model-free optimization approach can successfully identify non-intuitive, multi-faceted actuation strategies that yield significant and energetically efficient drag reduction.
@@ 1 Introduction
Road freight is responsible for approximately 5.2% of global greenhouse-gas emissions [itf_2021_report]. At highway speeds, roughly 65% of a vehicle’s tractive energy is expended to overcome aerodynamic drag [mccallen2004_aerodynamic_pollution_contribution]. For flat-backed vehicles such as vans and lorries, this drag is dominated by a low-pressure region on the base, which results from flow separation at the sharp trailing edges. Consequently, mitigating this base-pressure deficit through wake control provides a direct and effective means to reduce fuel consumption and the associated emissions.
@@
The wakes of simplified ground vehicles like the Ahmed body have been the subject of extensive research, revealing a highly complex, three-dimensional flow field that is acutely sensitive to geometric variations [Grandemange2013WakeBluntBody]. The near-wake is typically characterized by two large recirculation regions originating from the shear layers that separate from the upper and lower base edges [cerutti2020VanCafiero]. In the lateral direction, the wake often exhibits a global symmetry breaking, with modal switching between quasi-symmetric and anti-symmetric states. This low-frequency dynamic, linked to the shedding of large-scale vortex structures, produces broadband pressure fluctuations and a depressed time-averaged base pressure. Foundational experimental and numerical studies have thoroughly mapped the bi-modal character of these wakes and their sensitivity to geometry and operating conditions [dalla2019simulationsBimodalFlatBackBluff, fan2020experimentalBiStableFlatBackAhmed, khan2024equilibriumFluxesFlatBackBluffBody, Grandemange2013BistabilityBackAR], providing a baseline understanding that informs all modern control strategies.
@@
Modifying these wake dynamics offers a promising route for drag reduction. Passive control devices have been widely explored, including vertical flaps to attenuate lateral vortex structures [Beaudoin2008flaps], base cavities to improve performance in crosswinds [urquhart2021drag], splitter plates [gillieron2010aerodynamic], boat tails [lanser1991aerodynamic, khalighi2001experimental], and fences [modi1995drag]. In contrast, active flow control provides greater authority due to its adaptive capabilities to changing conditions [yu2021recentAdvances] and has been investigated extensively through momentum injection at the trailing edges [mcnally2015drag]. In particular, strategies involving steady momentum injection have proven effective, with roumeas2009drag achieving a 17% drag reduction on an Ahmed body via suction prior to separation, and aubrun2011separation obtaining up to 14% drag reduction using an array of steady blowing microjets. A widespread and often more efficient alternative is unsteady forcing, through pulsed or synthetic jets. These actuators can promote flow reattachment, as explored by glezer2005aspects, or directly target wake instabilities. Examples of the latter include parametric studies to optimize actuation on the slanted surfaces of an Ahmed body [park2013aerodynamic, joseph2013flow], and targeted base pressure recuperation using high-frequency forcing [oxlade2015high]. Some approaches have even combined passive devices with active suction and blowing to maximize performance [seifert2015lab].
@@
More recently, learning-based, model-free approaches have gained traction, offering powerful tools to navigate the vast and complex parameter space of active flow control laws directly on experimental setups. This paradigm bypasses the need for explicit, often intractable, low-dimensional models of turbulent wakes, allowing algorithms to discover effective, sometimes non-intuitive, control strategies through interaction with the physical system [li2017GP_FlatBackAhmed, Zhang2023AIAhmed, zhang2025autonomousrealtimecontrolturbulent]. Several successful applications have demonstrated the potential of these methods on benchmark automotive geometries, primarily the Ahmed body. For instance, li2017GP_FlatBackAhmed employed linear genetic programming (LGP) to optimize pulsed jets combined with Coanda surfaces on a square-back Ahmed body, achieving a remarkable 22% drag reduction. Their model-free approach identified optimal multi-frequency forcing laws. Zhang2023AIAhmed utilized an Ant Colony Optimization (ACO) algorithm to control distributed steady microjet arrays on a low-drag Ahmed body (slant angle $\varphi=35^{\circ}$), reducing drag by 18% while explicitly incorporating control power input into the cost function to seek efficient solutions. deng2023sensitivity applied an Explorative Downhill Simplex Method (EDSM) to optimize independently operated pulsed microjets on a square-back Ahmed body, achieving 13% drag reduction. Their work particularly focused on sensitivity analysis for a large number of control parameters (up to 12) and demonstrated the potential for significantly enhanced control efficiency (up to 78%) with only a minor sacrifice in drag reduction. Demonstrating a fully autonomous approach, zhang2025autonomousrealtimecontrolturbulent developed REACT, a Reinforcement Learning framework deployed experimentally on a square-back Ahmed body using servo-actuated flaps. Learning directly in the wind tunnel from onboard pressure sensors, REACT achieved a 3.64% drag reduction with net energy savings, discovering that dynamically suppressing wake instabilities was the optimal strategy. Notably, their physics-informed training allowed the agent to generalize across different flow speeds without retraining. Collectively, these successes underscore the competitive performance and adaptability of machine learning control for manipulating complex wake flows on standard automotive shapes
@@
Moving beyond the standard Ahmed body benchmark, active flow control strategies have also been applied to more realistic vehicle shapes, notably the van model geometry utilized in [amico2022geneticFurgo, amico2022deep, amico2024flow]. In a first study [amico2022geneticFurgo], a Genetic Algorithm (GA) was used to optimize multi-frequency pulsed jets in a square-back van (represented as a sum of two sine waves), achieving an 11.2% drag reduction. Notably, their initial cost function aimed purely at maximizing drag reduction without explicitly penalizing actuation energy. The follow-up work [amico2024flow] employed Deep Reinforcement Learning (DRL), exploring different state representations (using base pressure) and reward definitions, including cases explicitly designed to consider the energy budget alongside drag reduction. Their subsequent flow topology analysis revealed distinct wake modifications depending on whether the DRL agent prioritized maximal drag reduction or energy efficiency. The efficiency and adaptability of DRL have made it one of the most attractive closed-loop control methodologies not only for bluff-body flows, but more broadly within fluid mechanics. Representative examples include closed-loop attenuation of a cylinder wake using plasma actuators [Zong2024airfoilplasmaclosedloop], airfoil separation control [Zong2024airfoilplasmaclosedloop], and highly complex configurations such as the maximization of shear-layer mixing in a supersonic backward-facing step [Zong2025closedloopsupersonic].
@@
Despite the demonstrated promise, the efficacy and practical applicability of these learning-based strategies can be constrained by several factors. Firstly, the reliance on a single optimization algorithm (be it GA, LGP, ACO, EDSM, or DRL) within many individual studies risks incomplete exploration of the potentially vast and multimodal control parameter space, possibly leading to convergence towards locally optimal, rather than globally optimal, solutions. Secondly, a critical consideration, often omitted in earlier works but highlighted in more recent ones, is the actuation cost. Neglecting the energy expenditure of the control system within the optimization loop can yield strategies that, while achieving substantial drag reduction, are energetically inefficient or entirely impractical for real-world applications. While several studies have begun incorporating energy considerations [e.g. Zhang2023AIAhmed, deng2023sensitivity, amico2024flow], it remains a crucial factor for practical viability. Finally, many experimental implementations, particularly those involving hardware-in-the-loop optimization, do not formally account for the propagation of measurement uncertainty. The inherent noise in experimental data, if not properly managed, can be amplified, introducing significant bias into the learning process and potentially guiding the optimization towards spurious or non-robust outcomes.
@@
To address these limitations, the present study introduces a novel hybrid genetic algorithm that synergistically combines the global exploratory power of a genetic algorithm with the local exploitation capabilities of the Downhill Simplex Method, a local search algorithm. We apply this algorithm experimentally to a simplified van model in a wind tunnel as in amico2022geneticFurgo, amico2022deep, amico2024flow, optimizing the control parameters of a pulsed-jet system while simultaneously characterizing the resulting wake modifications. Crucially, the objective function incorporates a penalty term for the momentum injected by the actuators, ensuring that the optimization converges towards energetically feasible solutions. Furthermore, to guarantee the robustness of the experimental evaluation, each candidate solution is subjected to repeated measurements, and a statistical uncertainty threshold is employed to discard unreliable data, thereby providing a high degree of confidence in the final performance metrics.
@@
The remainder of this paper is organized as follows. Section 2 details the experimental setup, including the wind tunnel facility, the bluff body model, and the actuation and measurement systems. Section 3 presents the hybrid genetic algorithm, defining the control parameterization, the cost function, and the methodology for handling experimental uncertainty. The results of the optimization are presented and discussed in Section 4, where the best-performing control strategies is analyzed in detail, and the corresponding flow physics are examined in Section 4.3. Finally, Section 5 summarizes the key findings of the study and offers concluding remarks.
@@ 2.1 Experimental Facility and Model Setup
The experiments were conducted in an Eiffel-type wind tunnel. The facility has a test section with a square cross-section of $0.4\times 0.4$ m2 and a length of $2$ m. All measurements were performed at a free-stream velocity of $U_{\infty}=12$ m/s, with a corresponding turbulence intensity below 0.3%.
The bluff body is a scaled version of the van model used in cerutti2020VanCafiero and amico2022geneticFurgo, amico2022deep, amico2024flow, with a height of $H=0.12$ m, a width of $W=0.102$ m, and a length of $L=0.247$ m. The clearance between the model and the wind tunnel floor is equal to $h=0.012$ m, leaving a normalized value $h/H=0.1$. This configuration results in a wind tunnel blockage ratio of 7.65% and a Reynolds number based on width of $Re_{W}\approx 78,300$. It is noted that, even though the external geometry of the model is identical to the referenced study, the interface with the load transducer and the actuation systems is not the same (see §2.2). As depicted in Figure 1, the model is positioned on a free-floating platform, maintaining a lateral clearance of approximately $1$ mm with respect to the wind tunnel floor. It is rigidly mounted to the platform through rectangular extensions from the wheels, designed so that the wheel base would be flush with the floor. This mounting strategy provides a non-intrusive passage for the tubing of the actuation and instrumentation systems, thereby precluding the need for internal support structures and the associated flow corrections.
@@ 2.2 Actuation System
The actuation system comprises four pairs of parallel slots positioned on the four edges of the model’s base, as indicated by the red lines in Figure 1, (b). This arrangement is analogous to actuator configurations used in recent bluff-body flow-control studies, such as the single-slot jets employed by amico2022geneticFurgo. Each rectangular slot features a cross-section of $39.4\times 1$ mm2. The use of slot pairs, separated by $5$ mm on each side of the model, is employed to improve the homogeneity of the jet outflow.
@@ 2.3 Force and base Pressure Measurements
Aerodynamic drag force is measured via a custom lever system acting on a Fibos FA702 three-axis load cell. The body-platform assembly, described in Section 2.1, is mounted on a primary cantilever beam of length $L_{B}=0.82$ m. This beam pivots on a high-precision bearing, a mechanism that translates the streamwise aerodynamic force exerted on the model into a vertical load on a secondary lever arm of length $L_{b}=85$ mm. The resulting force is transmitted via a vertical threaded rod to the load cell. In the primary measurement axis, the load cell has a maximum capacity of $5$ N and sensitivity 1mV/V1\;mV/V with an excitation of 12V12\;V. Data from the load cell is acquired using a Viking VK702NH data acquisition system, which provides a $10$ V excitation voltage and utilizes a 24-bit ADC with an input range of $100$ mV, sampling at $800$ Hz. Finally, solid blockage [barlow1999low] and wake blockage correction using Maskell’s methodology [Gould1969WakeBlockage] is applied to the measured drag coefficients.